# Q-Bench-Video: Benchmarking the Video Quality Understanding of LMMs

### v1

@@ 3.3.4 Questions & Answers Annotation
The annotation process of Q-Bench-Video is conducted in a well-controlled laboratory environment. A total of 8 experts are employed and trained to ensure the consistency of the annotations. The experts are required to watch the videos in their entirety before making annotations. Each annotated question-answer pair is then reviewed by at least three other experts to ensure its validity and accuracy. The annotation details and GUI visualization are presented in Appendix C.
@@ 3.4 Benchmark Setting & Evaluation
Unless specifically stated otherwise, for Video LMMs we typically analyze by uniformly sampling 16 frames from the video, while for Image LMMs, the sampling is reduced to 8 frames. For Yes-or-No and What-How questions, if the LMMs can accurately respond with the options, we directly record the accuracy of the responses as results. If the LMMs cannot provide option-based answers, we implement a GPT-assisted evaluation strategy to help judge the accuracy of the answers. For Open-ended questions, since the answers are open-ended and cannot be directly quantified for accuracy, we also employ the GPT-assisted evaluation strategy. This involves GPT scoring the responses based on their accuracy, completeness, and relevance compared to the annotated answer. Details about benchmark setting and evaluation can be found in the Appendix E and Appendix G.
(a) Overall performance on Q-Bench-Video.
(b) Subcategory performance on Q-Bench-Video.
4 Results of Q-Bench-Video
4.1 Experimental Setting
LMMs Participants. A total of 17 LMMs (12 Open-source LMMs and 5 Proprietary LMMs) are included for validation, which includes a) 3 Open-source Image LMMs: LLaVA-Next (Liu et al., 2024a), LLaVA-v1.5 (Liu et al., 2023a), and mPLUG-Owl2 (Ye et al., 2023b); b) 9 Open-source Video LMMs: mPLUG-Owl3 (Ye et al., 2024), LLaVA-OneVision (Li et al., 2024a), InternVL-Chat (Chen et al., 2024), VILA1.5 (Ke et al., 2023), PLLaVA (Xu et al., 2024), LLaVA-Next-Video (Zhang et al., 2024a), ST-LLM (Liu et al., 2023c), Video-LLaVA (Lin et al., 2023), and VideoChat2 (Li et al., 2023a); c) 5 Proprietary LMMs: Gemini 1.5 Flash, Gemini 1.5 Pro (Team, 2024), GPT-4o mini, GPT-4o, and GPT-4 Turbo (Achiam et al., 2023).
Subsets Split. The Q-Bench-Video is divided into test (1,186 question-answer items) and dev (1,192 question-answer items) subsets. The correct answers will be released and proprietary for the dev and test subsets respectively. All discussions and analyses are based on the test subset. The details for the human performance on the test subset can be found in Appendix F.
Table 2: Results on the test subset for the video quality perception ability of LMMs. The best performance is marked in bold and the second performance is underlined for Open-source and Proprietary LMMs respectively. The Open-ended questions are judged as 0.00% for Random guess.
@@ 4.1 Experimental Setting
Random guess 50.00% 25.00% 0.00% 23.70% 23.46% 25.83% 21.69% 25.67%
@@ 4.2 Findings
The overall performance and subcategory comparisons (human vs. top-performing LMMs) on Q-Bench-Video can be quickly glanced at Fig. 3. Detailed performance across the subcategories for each LMM are shown in Table 2 (Question Types & Quality Concerns) and Table 3 (Single Videos vs. Video Pairs) respectively. The organization of the findings is as follows:
1) General Performance. Human>Proprietary LMMs>Open-source LMMs>Random guess. From the performance results presented in Table 2, we observe that nearly all LMMs significantly outperform random guess, demonstrating their basic capability to understand video quality. Among the open-source LMMs, the recently released mPLUG-Owl3 achieves the highest overall performance at 52.39%, even slightly surpassing GPT-4o mini (52.20%), followed closely by LLaVA-OneVision (51.70%) and InternVL-Chat (51.11%). Image LMMs deliver moderate performance. Although they outperform some Video LMMs, the gap between them and the latest Video LMMs is still notable. Benefiting from larger training datasets and more parameters, proprietary LMMs (except GPT-4o mini) outperform all open-source models. However, even the best-performing model, GPT-4o, which achieved an overall performance of 58.70%, still lags behind human performance by 22.86%. This gap highlights that, despite the advancements in current state-of-the-art LMMs, there remains a significant need for improvement in video quality understanding ability.
2) Question Types. Open-ended questions are more challenging for LMMs. From Table 2, a discernible hierarchy in task difficulty for video quality assessment emerges for both humans and LMMs, arranged as follows: Open-ended >What-How >Yes-or-No. It is crucial to highlight that while humans exhibit a performance decline in Open-ended tasks by approximately 9.46% compared to Yes-or-No tasks, and about 3.89% compared to What-How tasks, these reductions are markedly less pronounced than those observed in LMMs for the Open-ended questions. This disparity underscores a significant proficiency gap between LMMs’ capability in handling straightforward, closed-form questions and their effectiveness in navigating the complexities of real-world problem-solving, particularly in the context of video quality evaluation.
@@
Random guess 21.49% 27.08% 24.47% 29.58% 31.93% 27.40% 29.46%
@@ 5 Conclusion
In this paper, we introduce Q-Bench-Video, the first comprehensive benchmark explicitly designed to evaluate Large Multi-modal Models’ (LMMs) understanding of video quality. Our benchmark includes a diverse range of video types, questions that challenge multiple aspects of video quality, and a holistic evaluation framework encompassing Technical, Aesthetic, Temporal, and AIGC distortions. Through extensive experimentation with 17 open-source and proprietary LMMs, we find that while LMMs show promise in discerning video quality, their performance remains significantly below human-level understanding, especially when addressing Open-ended questions and AIGC-specific distortions. These findings highlight the current limitations of LMMs in video quality perception and underscore the need for further advancements in this area. By offering Q-Bench-Video, we aim to stimulate future research and drive improvements in the field, ultimately bridging the gap between LMM and human video quality assessment capabilities.
Appendix A Source Video Dataset Introduction
@@
  • LSVQ (Ying et al., 2021): The LSVQ dataset is currently the largest VQA dataset, comprising over 39,000 real-world videos and 5.5 million human perceptual quality annotations. It primarily focuses on both spatial and temporal aspects of technical visual quality.
  • MaxWell (Wu et al., 2023b): The MaxWell dataset presents a comprehensive subjective study, gathering over two million human opinions on 13 distinct quality factors across 4,543 in-the-wild natural scene videos. These quality factors include technical aspects (sharpness, focus, noise, motion blur, flicker, exposure, compression artifacts, fluency) as well as aesthetic aspects (content appeal, composition, color, lighting, and camera trajectory).
  • WaterlooSQoE-III (Duanmu et al., 2018): The WaterlooSQoE-III dataset comprises 20 RAW HD reference videos and 450 simulated streaming videos. To generate meaningful and representative test videos, a series of DASH video streaming experiments are conducted, capturing relevant streaming activities and reconstructing the streaming sessions using video processing tools. The WaterlooSQoE-III dataset primarily focuses on assessing the quality of experience (QoE) in streaming video.
  • WaterlooSQoE-IV (Duanmu et al., 2020): The WaterlooSQoE-IV dataset is currently the largest subject-rated VQA dataset for quality of experience, featuring 1,350 adaptive streaming videos. These videos are generated from a diverse range of source content, video encoders, network conditions, adaptive bitrate (ABR) algorithms, and viewing devices.
  • T2VQA-DB (Kou et al., 2024): The T2VQA-DB dataset utilizes 9 different video generation models to create 10,000 AIGC videos. A total of 27 subjects are invited to assess the perceptual quality of each video, focusing on two main aspects: text-video alignment and video fidelity. Text-video alignment refers to how well the generated video content corresponds to the given text description, while video fidelity encompasses factors such as distortion, saturation, motion consistency, and content coherence.
  • VideoFeedback (He et al., 2024b): The VideoFeedback dataset contains human-provided multi-aspect scores for 37.6K synthesized videos generated by 11 different video generative models. The scores assess various aspects, including visual quality, temporal consistency, dynamic realism, text-to-video alignment, and factual consistency.
  • LIVE-YT-Gaming (Yu et al., 2023): The LIVE-YT-Gaming dataset consists of 600 real user-generated gaming videos. A subjective human study is conducted on this dataset, resulting in 18,600 quality ratings provided by 61 participants. The primary focus of the study is on evaluating the visual quality of the videos.
@@
Appendix B Video Sampling Approach
@@
$[q_{\text{min}},q_{\text{min}}+\Delta q),\quad[q_{\text{min}}+\Delta q,q_{\text{min}}+2\Delta q),\quad\dots,\quad[q_{\text{min}}+4\Delta q,q_{\text{max}}],$ (3)
@@
(a) Annotation GUI for single videos of Q-Bench-Video.
(b) Annotation GUI for video pairs of Q-Bench-Video.
@@
Appendix C Annotation Process
A group of eight experts, all with professional photography skills and extensive experience, participate in the subjective annotation experiment for Q-Bench-Video. The experiment takes place in a controlled lab environment with standard indoor lighting. A Dell 4K monitor with a resolution of 3840 × 2160 is used to display the interfaces, as shown in Fig 4. To avoid fatigue, each expert labels up to 30 videos per day, and every annotation is carefully reviewed by at least three other experts before final approval. This process ensures the highest level of accuracy and rigor in the Q-Bench-Video labels, making performance testing of Q-Bench-Video more precise and meaningful.
Appendix D Yes-or-No Ratio
As indicated in numerous previous works (Zhang et al., 2024c; Wu et al., 2024b), LMMs often exhibit a bias when answering Yes-or-No questions, such as a tendency to favor Yes over No. To mitigate this issue, we specifically examine the distribution of correct answers for Yes-or-No questions and adjust the questions to ensure a balanced 50%/50% ratio between Yes and No answers. For illustration, we provide an example to demonstrate how we modify the questions:
@@ Appendix D Yes-or-No Ratio
Appendix E Benchmark Setting
E.1 Sampling Frames.
@@ Appendix E Benchmark Setting
E.2 Prompt for single videos on MCQ
@@
E.3 Prompt for single videos on Open-ended Questions
@@
E.4 Prompt for video pairs on MCQ
@@
E.5 Prompt for video pairs on Open-ended Questions
@@
Appendix F Human Performance on Q-Bench-Video
To evaluate the gap between LMMs and human performance in video quality understanding, we invite three human participants to take part in experiments using the test subset of Q-Bench-Video. The experimental setup and procedure are identical to the annotation environment previously described (See Appendix C). It’s worth noting that the participants undergo a brief training session to familiarize themselves with the tasks and acquire the necessary knowledge of video quality. Afterward, they complete the test, and we record their average scores as the final results of human performance. The human performance testing interface is shown in Fig. 5.
Figure 6: Qualitative comparison for LMM on MCQ response.
Appendix G Evaluation Details
G.1 Evaluation on MCQs
@@
G.2 MCQ Evaluation Example
@@
G.3 Evaluation on Open-ended Questions
$T=\sum_{i=1}^{5}S_{i}$
$\text{Accuracy}=\frac{T}{10}$
@@ G.3 Evaluation on Open-ended Questions
Random guess 50.00% 25.00% 0.00% 25.74% 21.98% 26.56% 25.54% 27.14%
@@
G.4 Open-ended Question Evaluation Example
5-round GPT score: [0, 0, 1, 0, 0] Final Score: 1/10 = 0.1
Appendix H Qualitative Visualization Results
In Fig. 6, we present an example of LMM responses to MCQ questions. First, it is clear that most LMMs can follow the instructions well to select the option they believe is correct, except for ST-LLM, which requires further checking to determine if the answer is correct (See Appendix G for details about GPT-assisted evaluation approach). Secondly, for basic quality assessments like clarity, which are relatively simple for humans, about half of the LMMs still failed to answer correctly. This again highlights the need to improve LMM video understanding capabilities.
Appendix I Performance on the dev subset

### v2

@@ 3.3.4 Questions & Answers Annotation
The annotation process of Q-Bench-Video is conducted in a well-controlled laboratory environment. A total of 8 experts are employed and trained to ensure the consistency of the annotations. The experts are required to watch the videos in their entirety before making annotations. Each annotated question-answer pair is then reviewed by at least three other experts to ensure its validity and accuracy.
@@ 3.4 Benchmark Setting & Evaluation
Unless specifically stated otherwise, for Video LMMs we typically analyze by uniformly sampling 16 frames from the video, while for Image LMMs, the sampling is reduced to 8 frames. It’s worth noting that if the LMMs are not compatible with the specified number of sampling frames (e.g. if they cannot support up to 8-frame input), we will use the official default settings for testing. For Yes-or-No and What-How questions, if the LMMs can accurately respond with the options, we directly record the accuracy of the responses as results. If the LMMs cannot provide option-based answers, we implement a GPT-assisted evaluation strategy to help judge the accuracy of the answers. For Open-ended questions, since the answers are open-ended and cannot be directly quantified for accuracy, we also employ the GPT-assisted evaluation strategy. This involves GPT scoring the responses based on their accuracy, completeness, and relevance compared to the annotated answer.
Table 2: Results on the test subset for the video quality perception ability of LMMs. The best performance is marked in bold and the second performance is underlined for Open-source and Proprietary LMMs respectively. The Open-ended questions are excluded for Random guess. (Performance of the test subset is in the Supp.)
@@
     Random guess w/o Open-ended      50.00%      25.00%      /      37.10%      37.31%      37.25%      37.22%      37.79%
@@
4 Results of Q-Bench-Video
4.1 Experimental Setting
LMMs Participants. A total of 17 LMMs (12 Open-source LMMs and 5 Proprietary LMMs) are included for validation, which includes a) 3 Open-source Image LMMs: LLaVA-Next [30], LLaVA-v1.5 [28], and mPLUG-Owl2 [54]; b) 9 Open-source Video LMMs: mPLUG-Owl3 [52], LLaVA-OneVision [19], InternVL-Chat [4], VILA1.5 [16], PLLaVA [50], LLaVA-Next-Video [58], ST-LLM [31], Video-LLaVA [27], and VideoChat2 [22]; c) 5 Proprietary LMMs: Gemini 1.5 Flash, Gemini 1.5 Pro [39], GPT-4o mini, GPT-4o, and GPT-4 Turbo [2].
Subsets Split. The Q-Bench-Video is divided into test (1,186 question-answer items) and dev (1,192 question-answer items) subsets. The correct answers will be released and proprietary for the dev and test subsets. All discussions and analyses are based on the test subset.
@@ 4.2 Findings
The overall performance and subcategory comparisons (human vs. top-performing LMMs) on Q-Bench-Video can be quickly glanced at Fig. 4. Detailed performance across the subcategories for each LMM are shown in Table 2 (Question Types & Quality Concerns) and Table 3 (Single Videos vs. Video Pairs) respectively. The findings are organized as follows:
1) General Performance. Human>Proprietary LMMs>Open-source LMMs>Random guess. From the performance results presented in Table 2, we observe that nearly all LMMs significantly outperform random guess, demonstrating their basic capability to understand video quality. Among the open-source LMMs, the recently released mPLUG-Owl3 achieves the highest overall performance at 52.39%, even slightly surpassing GPT-4o mini (52.20%), followed closely by LLaVA-OneVision (51.70%) and InternVL-Chat (51.11%). Image LMMs deliver moderate performance. Although they outperform some Video LMMs, the gap between them and the latest Video LMMs is still notable. Benefiting from larger training datasets and more parameters, proprietary LMMs (except GPT-4o mini) outperform all open-source models. However, even the best-performing model, GPT-4o, which achieved an overall performance of 58.70%, still lags behind human performance by 22.86%. This gap highlights that, despite the advancements in current state-of-the-art LMMs on video general understanding, there remains a significant need for improvement in video quality understanding ability.
@@
     Random guess w/o Open-ended      36.49%      38.61%      37.71%      40.56%      37.94%      36.83%      38.00%
@@
2) Question Types. Open-ended questions are more challenging for LMMs. From Table 2, a discernible hierarchy in task difficulty for video quality assessment emerges for both humans and LMMs, arranged as follows: Open-ended >What-How >Yes-or-No. It is crucial to highlight that while humans exhibit a performance decline in Open-ended tasks by approximately 9.46% compared to Yes-or-No tasks, and about 3.89% compared to What-How tasks, these reductions are markedly less pronounced than those observed in LMMs for the Open-ended questions. This disparity underscores a significant proficiency gap between LMMs’ capability in handling straightforward, closed-form questions and their effectiveness in navigating the complexities of real-world problem-solving on open-ended questions, particularly in the context of video quality evaluation.
5 Limitations
1) Subjectivity can not be fully avoided in annotations for aspects like aesthetics. 2) As video generation advances, new models may show different distortions, which might potentially make parts of Q-Bench-Video outdated.
@@ 6 Conclusion
In this paper, we introduce Q-Bench-Video, the first comprehensive benchmark explicitly designed to evaluate Large Multi-modal Models’ (LMMs) understanding of video quality. Our benchmark includes a diverse range of video types, questions that challenge multiple aspects of video quality, and a holistic evaluation framework encompassing Technical, Aesthetic, Temporal, and AIGC distortions. Through extensive experimentation with 17 open-source and proprietary LMMs, we find that while LMMs show promise in discerning video quality, their performance remains significantly below human-level understanding, especially when addressing Open-ended questions and AIGC-specific distortions. These findings highlight the current limitations of LMMs in video quality perception and underscore the need for further advancements in this area. By offering Q-Bench-Video, we aim to stimulate future research and drive improvements in the field, bridging the gap between LMM and human video quality understanding.
Supplementary Material
7 Source Video Dataset Introduction
@@
  • LSVQ [55]: The LSVQ dataset is currently the largest VQA dataset, comprising over 39,000 real-world videos and 5.5 million human perceptual quality annotations. It primarily focuses on both spatial and temporal aspects of technical visual quality.
  • MaxWell [46]: The MaxWell dataset presents a comprehensive subjective study, gathering over two million human opinions on 13 distinct quality factors across 4,543 in-the-wild natural scene videos. These quality factors include technical aspects (sharpness, focus, noise, motion blur, flicker, exposure, compression artifacts, fluency) as well as aesthetic aspects (content appeal, composition, color, lighting, and camera trajectory).
  • WaterlooSQoE-III [6]: The WaterlooSQoE-III dataset comprises 20 RAW HD reference videos and 450 simulated streaming videos. To generate meaningful and representative test videos, a series of DASH video streaming experiments are conducted, capturing relevant streaming activities and reconstructing the streaming sessions using video processing tools. The WaterlooSQoE-III dataset primarily focuses on assessing the quality of experience (QoE) in streaming video.
  • WaterlooSQoE-IV [7]: The WaterlooSQoE-IV dataset is currently the largest subject-rated VQA dataset for quality of experience, featuring 1,350 adaptive streaming videos. These videos are generated from a diverse range of source content, video encoders, network conditions, adaptive bitrate (ABR) algorithms, and viewing devices.
  • T2VQA-DB [17]: The T2VQA-DB dataset utilizes 9 different video generation models to create 10,000 AIGC videos. A total of 27 subjects are invited to assess the perceptual quality of each video, focusing on two main aspects: text-video alignment and video fidelity. Text-video alignment refers to how well the generated video content corresponds to the given text description, while video fidelity encompasses factors such as distortion, saturation, motion consistency, and content coherence.
  • VideoFeedback [13]: The VideoFeedback dataset contains human-provided multi-aspect scores for 37.6K synthesized videos generated by 11 different video generative models. The scores assess various aspects, including visual quality, temporal consistency, dynamic realism, text-to-video alignment, and factual consistency.
  • LIVE-YT-Gaming [56]: The LIVE-YT-Gaming dataset consists of 600 real user-generated gaming videos. A subjective human study is conducted on this dataset, resulting in 18,600 quality ratings provided by 61 participants. The primary focus of the study is on evaluating the visual quality of the videos.
@@
8 Video Sampling Approach
@@
, (3)
@@
((a))
((b))
@@
9 Annotation Process
A group of eight experts, all with professional photography skills and extensive experience, participate in the subjective annotation experiment for Q-Bench-Video. The experiment takes place in a controlled lab environment with standard indoor lighting. A Dell 4K monitor with a resolution of 3840 × 2160 is used to display the interfaces, as shown in Fig 3. To avoid fatigue, each expert labels up to 30 videos per day, and every annotation is carefully reviewed by at least three other experts before final approval. Specifically, the raters statistics are [X (number of raters with agreement):(proportion)]: {[4, 85.1%] (which we only use), [3, 10.2%], [2, 4.7%]}. This process ensures the highest level of accuracy and rigor in the Q-Bench-Video labels, making performance testing of Q-Bench-Video more precise and meaningful.
10 Yes-or-No Ratio
As indicated in numerous previous works [62, 47], LMMs often exhibit a bias when answering Yes-or-No questions, such as a tendency to favor Yes over No. To mitigate this issue, we specifically examine the distribution of correct answers for Yes-or-No questions and adjust the questions to ensure a balanced 50%:50% ratio between Yes and No answers. For illustration, we provide an example to demonstrate how we modify the questions:
@@ 10 Yes-or-No Ratio
Table 4: Performance results of uniform 16 frames extraction and 1 FPS extraction evaluation settings, shown as (Uniform16/1FPS).
[…]
mPLUG3-7B (0.526/0.501) (0.583/0.581) (0.520/0.544) (0.434/0.447) (0.523/0.525)
InternVL-7B (0.484/0.472) (0.527/0.531) (0.505/0.510) (0.531/0.542) (0.511/0.520)
LLaVA-OV-7B (0.493/0.512) (0.641/0.655) (0.506/0.513) (0.443/0.440) (0.517/0.528)
Gemini 1.5 Pro (0.568/0.563) (0.696/0.687) (0.533/0.523) (0.532/0.540) (0.586/0.580)
GPT-4o (0.560/0.562) (0.585/0.591) (0.653/0.661) (0.522/0.513) (0.587/0.592)
11 Benchmark Setting
11.1 Sampling Frames.
@@ 11 Benchmark Setting
11.2 Prompt for single videos on MCQ
@@
11.3 Prompt for single videos on Open-ended Questions
@@
11.4 Prompt for video pairs on MCQ
@@
11.5 Prompt for video pairs on Open-ended Questions
@@
11.6 1 FPS evaluation setting:
We further consider the 1 FPS setting in our benchmark/ Given that most videos in the benchmark are between 10-16s, the gap between uniformly sampling 16 frames and 1 FPS frame extraction is small for our benchmark. We have tested some LMMs under this setting, with results shown in Table 4. The performance gap between these two settings is slight for our videos in general.
@@
12 Human Performance on Q-Bench-Video
To evaluate the gap between LMMs and human performance in video quality understanding, we invite three human participants to take part in experiments using the test subset of Q-Bench-Video. The experimental setup and procedure are identical to the annotation environment previously described (See Appendix 9). It’s worth noting that the participants undergo a brief training session to familiarize themselves with the tasks and acquire the necessary knowledge of video quality. Afterward, they complete the test, and we record their average scores as the final results of human performance. The human performance testing interface is shown in Fig. 6.
13 Evaluation Details
13.1 Evaluation on MCQs
@@
Figure 7: Qualitative comparison for LMM on MCQ response.
13.2 MCQ Evaluation Example
@@
13.3 Open-sourced LLM as Judge
Considering that GPT-4o is not stable and may change over time, we provide results using the open-source LLM Qwen2.5-72B [51] as the judge model for reference. The judge results are similar to GPT-4o (shown in Table 5) but offer greater consistency since open-source LLMs are fixed.
Table 5: Open-ended performance of using Qwen2.5-72B and GPT as judges. The rank order for these LMMs is unchanged.
Judge mPLUG3-7B InternVL-7B VLLaVA-7B Gemini 1.5 pro GPT-4o
Qwen 0.379/rank3 0.311/rank4 0.302/rank5 0.460/rank2 0.489/rank1
GPT 0.394/rank3 0.339/rank4 0.291/rank5 0.476/rank2 0.492/rank1
@@ 13.3 Open-sourced LLM as Judge
Random guess w/o Open-ended 50.00% 25.00% / 36.76% 37.30% 37.27% 37.62% 37.27%
@@
13.4 Evaluation on Open-ended Questions
$T=\sum_{i=1}^{5}S_{i}$
$\text{Accuracy}=\frac{T}{10}$